OpenStudy (anonymous):

[06.03] 15.0g of an unknown substance is heated from 25.0°C to 34.0°C. If 60.6 J of energy is absorbed, identify the unknown substance using the specific heat.

OpenStudy (anonymous):

Q=m*c*∆t 60.6=15*c*(34-25) Solve for c, then google or check the textbook for the substance with the specific heat capacity!